OpenStudy (anonymous):

Find the first six partial sums S1, S2, S3, S4, S5, S6 of the sequence. 4, 7, 10, 13, . . . can someone show me the steps?

OpenStudy (anonymous):

If the sequence is \(\{a_n\}\), then the nth partial sum is given by \[s_n = a_1 + a_2 + \cdots + a_{n-1} + a_n.\] Therefore, in this case, \(s_1 = 4\), \(s_2 = 4+7\), \(s_3 = 4+7+10\), etc.

OpenStudy (anonymous):

thanks